# Polynomial-like dynamics of analytic maps

### v1

@@ 1.3. Main results
Corollary 1.1.
For the map $g:U^{\prime}\to U$, there exists a unique maximal (by inclusion) full completely invariant compact $K_{*}\subset U$ and one and only one of the following alternatives holds:
  1. (1)
    $K_{*}=\emptyset$,
  2. (2)
    $P_{g}\subset K_{*}$ and $K_{*}$ is the filled Julia set of a PL restriction as in Theorem 1.1; moreover, given $x\in U\setminus K_{*}$, either $g^{n}(x)\in U\setminus U^{\prime}$ for some $n\geq 0$, or the forward orbit $O(x)$ is well-defined but
    $\omega(x)\subset\partial U,$
  3. (3)
    $\partial K_{*}\subset P_{g}$ and $C_{g}\setminus K_{*}\neq\emptyset$; in fact, $K_{*}=\{z:GO(z)\subset\hat{P}_{g}\}$,
Furthermore, if $Z$ is any completely invariant compact set (not necessarily full) of $g:U^{\prime}\to U$, then $Z\subset K_{*}$.
Additionally, let $R_{g}$ be the closure (in the Riemann sphere $\hat{\mathbb{C}}$) of the set of repelling periodic points of $g:U^{\prime}\to U$. Then $R_{g}=\partial K_{*}$ in the case (2) while in cases (1) and (3) the compact $R_{g}$ (in $\hat{\mathbb{C}}$) contains a (non trivial) continuum which has common points with the range $U$ as well as with its boundary $\partial U$.
Note that $K_{*}\subset\cap_{i\geq 0}U_{-i}$ and $x\in K_{*}$ if and only if $GO(x)\subset K_{*}$. In particular, $C_{g}\setminus K_{*}\neq\emptyset$ holds if and only if $P_{g}\nsubseteq K_{*}$.
See Sect 5 for the proof of the corollary.
Remark 1.2.
It is not clear whether the alternative (3) can occur. Our proof of Theorem 1.1 uses the fact that $P_{g}\subset K$.
The proof of Theorem 1.1 is a mixture of dynamical and geometric arguments. First, we show that either $g:U^{\prime}\to U$ admits a PL restriction as in Theorem 1.1 or it has a lot of repelling periodic orbits close to $\partial U$, see Proposition 2.2 and Corollary 2.3. Therefore, Theorem 1.1 will be an immediately consequence of the next
Theorem 1.2.
Under the conditions of Theorem 1.1, the map $g:U^{\prime}\to U$ has no repelling periodic points outside of $K$.
The proof of Theorem 1.2 in the case of connected $K$ is rather straightforward: using the Riemann map $\pi:\{|z|>1\}\to\mathbb{C}\setminus K$ one can conjugate the map $g:U^{\prime}\setminus K\to U\setminus K$ to an unbranched $d$-covering map in a one-sided neighborhood of the unit circle $S^{1}$ and apply Schwartz’s reflection principle to extend it through $S^{1}$; then one can end the proof, for example, by noticing that a repelling cycle outside $S^{1}$ would attract inverse iterates of points of $S^{1}$, in contradiction with the complete invariance of $S^{1}$. It is not clear how to generalize this idea to the disconnected $K$. We propose another proof in the case of connected $K$, which is based on a different idea, see Sect. 3.1, and then carry this idea over to the disconnected case. The latter part is more technical and uses general results of the ergodic theory of locally analytic maps and the uniformization theory of plane sets, see Appendices A-B.
1.4. A generalization
Note that a priori the completely invariant compact $K$ of Theorem 1.1 is not required to be connected although a posteriori $K$ is connected because it contains (iterates of) all critical points of the PL map $g:V^{\prime}\to V$.
We have the following generalization of Theorem 1.1 where $K$ is always disconnected.
Remind that a generalized PL map [4] $G:V^{\prime}\to V$ is a holomorphic map between open sets $V^{\prime},V$, where $V$ is a topological disk, $\overline{V^{\prime}}\subset V$, and $V^{\prime}$ consists of a finitely many simply connected components $V_{k}^{\prime}$, k=1,,Mk=1,...,M, with pairwise disjoint closures, such that $G:V^{\prime}\to V$ is a proper map of some degree $D\geq 2$. Its set of non-escaping points $K_{G}=\{z:G^{n}(z)\in V^{\prime},n=0,1,$...\} - the filled Julia set of $G$ - is a completely invariant full compact. (Note that if $V^{\prime}$ is connected, $G:V^{\prime}\to V$ is the classical Douady-Hubbard PL map.) The Straightening theorem also holds: $G:V^{\prime}\to V$ is conjugated, in a neighborhood of $K_{G}$, to a polynomial of degree $D$ (see e.g. [3] for a proof of a similar claim). If $M\geq 2$, then $K_{G}$ is never connected.
Theorem 1.3.
Suppose that $U^{\prime}\subset U$ are open subsets of $\mathbb{C}$ where $U$ is connected simply connected and $U^{\prime}$ consists of a finitely many simply connected components $U^{\prime}_{k}$, k=1,,mk=1,...,m, for some $m\geq 2$. Let $g:U^{\prime}\to U$ be a proper holomorphic map of some degree $d\geq 2$. Let $K\subset U$ be a full compact set such that $g^{-1}(K)=K$ and $K$ contains all critical points of $g:U^{\prime}\to U$ (if any).
Then there exist a topological disk $V_{0}$, $\overline{V_{0}}\subset U$, and $N\geq 1$ such that $\overline{V_{-N}}\subset V_{0}$, where $V_{-N}=g^{-N}(V_{0})$, and $g^{N}:V_{-N}\to V_{0}$ is a generalized PL map of degree $d^{N}$ and $K$ is its filled Julia set.
Moreover, there exists a neighborhood $V\subset V_{0}$ of $K$ consisting of a finitely many simply connected components as follows: if $V^{\prime}:=g^{-1}(V)$, then $\overline{V^{\prime}}\subset V$ and $g:V^{\prime}\to V$ is a proper map of degree $d$, such that $K=\{x:g^{n}(x)\in V^{\prime},n=0,1,2,$...\}.
Remark 1.3.
Let again $g:U^{\prime}\to U$ be a proper holomorphic map of some degree $d\geq 2$ where $U^{\prime},U$ are simply connected domains, and $K$ is a full compact which is completely invariant. Suppose that not all critical points of $g$ are in $K$ and that any critical point $c$ of $g$ outside of $K$ escapes, i.e., $g^{n}(c)\in U\setminus U^{\prime}$ for some $n=n(c)>0$. Then Theorem 1.3 applies to some iterate $g^{M}:g^{-M}(U)\to U$ of $g$ (with e.g. $M=\max\{n(c)\}$) and the compact $K$.
Acknowledgments. I would like to thank Feliks Przytycki for several fruitful discussions concerning mainly Theorem A.1, and Misha Sodin, Sebastian van Strien and Vladlen Timorin for helpful comments and references.
2. Preliminary statements
Lemma 2.1.
Basin of every attracting or parabolic cycle of $g:U^{\prime}\to U$ contains a critical point of $g$.
Proof.
The proof repeats the classical one: assume that the immediate basin $B$ of some attracting or parabolic cycle $P$ of $g$ contains no critical values of $g$. Then the inverse of Konigs (if $P$ is attracting and not superattracting) or Fatou (if $P$ is parabolic) linearizing coordinate extends, by the inverse branches of $g$, to an entire function with values in $B$, a contradiction with Liouville’s theorem. ∎
@@ 2. Preliminary statements
Next, we are going to describe $3$ alternatives. Let us fix a Jordan curve $\Gamma_{0}$ in $U$ which encloses the postcritical set $P_{g}$. Let $V_{0}$ be the Jordan domain bounded by $\Gamma_{0}$, so that $P_{g}\subset V_{0}\subset\overline{V_{0}}\subset U$.
@@
Let, for each $n>0$, $V_{-n}=g^{-n}(V_{0})$ (which is compactly contained in $g^{-n}(U)$) and $\Gamma_{n}=\partial V_{-n}$ As $P_{g}\subset V_{0}$, then $V_{-n}$ is also simply connected and $\Gamma_{n}$ is a Jordan curve.
@@
Assume that (A) holds. Then, for every $\varepsilon>0$ and every $n$ big enough, the map $g^{n}$ has a repelling fixed point in $\varepsilon$-neighborhood of $\Gamma_{0}$.
Assume that (B) holds. Then $g^{n}:V_{-n}\to V_{0}$ is a PL map of degree $d^{n}$. Furthermore, if there exists a full compact $K\subset V_{0}$ such that $g^{-1}(K)=K$, then $K=K_{G}=\cap_{k\geq 0}g^{-kn}(V)$, the filled Julia set of the PL map $g^{n}:V_{-n}\to V_{0}$. Moreover, there exists a polynomial-like restriction $g:V^{\prime}\to V$ where $V$ is a simply connected neighborhood of $K$ and $K$ is the filled Julia set of $g:V^{\prime}\to V$. In particular, $P_{g}\subset K$
Assume that (C) holds. Then the restricted map $g:\tilde{V}^{\prime}\to\tilde{V}$ where $\tilde{V}^{\prime}=\cup_{i\geq 1}g^{-i}(V_{0})$, $\tilde{V}=\cup_{i\geq 0}g^{-i}(V_{0})$ is trivial: there is an attracting fixed point $a\in\cap_{i\geq 0}g^{-i}(V_{0})$ such that $g^{n}(z)\to a$ uniformly in $V$.
Moreover, if $\Gamma_{0}$ is close enough to $\partial U$, then (C) does not happen.
@@
(A). In this case, there exist a sequence of indices $n_{i}\to\infty$ and a sequences of points $x_{i}\in\Gamma_{n_{i}}\cap\Gamma_{0}$ as follows: $x_{i}\to b^{\prime}$, $y_{i}:=g^{n_{i}}(x_{i})\in\Gamma_{0}$, $y_{i}\to b$ and $(g^{n_{i}})^{\prime}(x_{i})\to L$ where $L\in\mathbb{C}\cup\{\infty\}$. Fix a small disk $B$ centered at $b$ which is contained in $U$ and disjoint with $P_{g}$, and, for each $i$, let $\phi_{i}$ be a branch of $g^{-n_{i}}$ which is well-defined in $B$ and maps $y_{i}$ to $x_{i}$. Then $(\phi_{i})^{\prime}(y_{i})\to 1/L$.
Thus $\phi_{i}:B\to\mathbb{C}$, $i=1,2,$..., is a sequence of univalent maps and is a normal family because it avoids taking values in the set $P_{g}\cup g^{-1}(P_{g})$, which must contain at least $2$ points. By this reason, the case $L=0$ is impossible.
Consider the case $L\neq\infty$, that is, $0<|L|<\infty$. Replacing $(n_{i})$ by a subsequence, in this case $(\phi_{i})$ tends uniformly on compacts in $B$ to a non-constant, hence, univalent function $\nu$ which maps $B$ to a neighborhood $B^{\prime}$ of $b^{\prime}$. As $\psi_{i}(B)\subset g^{-n_{i}}(U)$, then $B^{\prime}\subset\cap_{i\geq 0}g^{-i}(U)$. Shrinking slightly $B^{\prime}$, we get:
@@
Let $m_{i}=n_{n+1}-n_{n_{i}}$ Passing again to a subsequence of $(n_{i})$, one can assume that $m_{i}\to\infty$. Write $g^{n_{i+1}}=g^{m_{i}}\circ g^{n_{i}}$, on $B^{\prime}$. Here, $g^{n_{i}}(B^{\prime})\subset g^{n_{i}}(g^{-n_{i+1}}(U))\subset g^{-m_{i}}(U)$.
Shrinking again $B^{\prime}$ if necessary, by the Claim, $g^{m_{i}}\to\operatorname{Id}$, uniformly on $\nu^{-1}(B^{\prime})$ where $\nu^{-1}(B^{\prime})\subset g^{-m_{i}}(U)$. Concluding this argument (and shrinking $B$ to a smaller disk around $b$), we obtain that, for some branches $I_{i}:=g^{-m_{i}}$ on $B$, $I_{i}\to\operatorname{Id}$ uniformly on $B\subset U_{0}$.
Let now $D$ be any simply connected subdomain of $U\setminus P_{g}$ containing $B$. As all branches of $g^{-m_{i}}$ are well-defined in $D$ and form a normal family, passing to a subsequence, we get that $I_{i}$ (more precisely, their extensions from $B$ to $D$) converge uniformly on compacts in $D$ to $\operatorname{Id}$. As each $I_{i}$ is an inverse to $g^{m_{i}}$, it follows that, given a compactly contained in $D$ domain $F\subset D$, for all $i$ big enough, we have: $F\subset g^{-m_{i}}(U)$ and $g^{m_{i}}\to\operatorname{Id}$ uniformly on $F$.
We apply this to two domains $D_{1},D_{2}$ which together cover the curve $\Gamma_{0}$, and to their compactly contained subdomains $F_{1}\subset D_{1}$, $F_{2}\subset D_{2}$, such that $F_{1}\cup F_{2}\supset\Gamma_{0}$. As a result, we obtain that $\Gamma_{0}\subset\cap_{i\geq 0}g^{-m_{i}}(U)$ for all $i$ big enough and $g^{m_{i}}\to\operatorname{Id}$ uniformly on $\Gamma_{0}$. This is a contradiction by the following reason: on the one hand, as we have just shown, $\Gamma_{0}\subset g^{-m_{i}}(U)$, also $\Gamma_{0}=\partial V_{0}$ and $g^{m_{i}}(g^{-m_{i}}(U))=U$, hence, the sequence $g^{m_{i}}$ is well-defined on $\overline{V_{0}}$ and converges uniformly to the identity map $\operatorname{Id}$. Then $(g^{m_{i}})^{\prime}\to 1$ uniformly on compacts in $V_{0}$. On the other hand, $(g^{m_{i}})^{\prime}(c)=0$ for each $c\in C_{g}$ This contradiction excludes the case $0<|L|<\infty$.
@@
(B). $\overline{V}_{-n}\subset V_{0}$ for some fixed $n>0$. Then $G:=g^{n}:V_{-n}\to V_{0}$ is a polynomial-like map (of degree $d^{n}$). Let $K_{G}$ be the filled Julia set of $G$. Then $K_{G}=\cap_{i\geq 0}G^{-i}(V_{0})$.
Assume now that additionally there exists a compact $K$ as in the condition, i.e., $K\subset V_{0}$ and $K$ is full and completely invariant. This implies that $G^{-1}(K)=K$ and that $K\subset\cap_{i\geq 0}G^{-i}(V_{0})=K_{G}$. As $G$ is conjugate to a polynomial near $K_{G}$, it follows that $K=K_{G}$. Moreover, as $P_{G}=P_{g}\subset V_{0}$, hence, $K=K_{G}$ is connected.
@@
It remains to show $g$ itself has a PL restriction, i.e., there exist a simply connected neighborhood $V$ of $K$, such that $V^{\prime}:=g^{-1}(V)$ is compactly contained in $V$. If $n=1$, there is nothing to prove (set $V=V_{0}$). So let $n\geq 2$.
@@
We present two proofs: the first one uses heavily the fact that $K$ is a continuum while the second one holds for non-connected $K$ as well and can be easily adapted to the set up of Theorem 1.3 (where $K$ is disconnected).
First proof.
@@
The following argument is well-known. Let $\Psi:\mathbb{C}\setminus K\to\mathbb{D}^{*}:=\{|w|>1\}$ be the conformal isomorphism such that $\Psi(\infty)=\infty$. Below, $\hat{A}=\Psi(A\setminus K)$ for a neighborhood $A$ of $K$, and $A_{*}$ is the union of $\hat{A}$ with its symmetric w.r.t. $S^{1}$ completed by $S^{1}$. Similar notations are used for maps. Replace $g:U_{-1}\setminus K\to U_{0}\setminus K$ by $\hat{g}:=\Psi\circ g\circ\Psi^{-1}:\hat{U}_{-1}\to\hat{U}_{0}$. By Schwartz’s reflexion principle, $\hat{g}$ extends to a holomorphic map $g_{*}:(U_{-1})_{*}\to(U_{0})_{*}$. The latter map has no critical points as $g_{*}^{-1}(S^{1})=S^{1}$. Let $\hat{G}=\Psi\circ g\circ\Psi^{-1}:\hat{U}_{-m}\to\hat{U}_{0}$ and $G_{*}:(U_{-m})_{*}\to(U_{0})_{*}$. As this map is a covering and $\overline{(U_{-m})_{*}}\subset(U_{0})_{*}$, then $G_{*}$ expands the hyperbolic metric $\rho$ of the domain $(U_{0})_{*}$, i.e, $D_{G_{*}}(w):=\rho\circ G_{*}(w)|G_{*}^{\prime}(w)|/\rho(w)>1$ for $w\in(U_{-m})_{*}$. Now note that $G_{*}=g_{*}^{m}$ in some neighborhood of $S^{1}$. Define a new metric $\tilde{\rho}=\sum_{i=0}^{m-1}\rho\circ g_{*}^{i}|(g_{*}^{i})^{\prime}|$. Then $\tilde{D}_{g_{*}}(w):=\frac{\tilde{\rho}\circ g_{*}(w)|g_{*}^{\prime}(w)|}{\tilde{\rho}(w)}=1+\frac{\rho(w)}{\tilde{\rho}(w)}\{D_{G_{*}}(w)-1\}>1$ for all $w$ near $S^{1}$. By the compactness, $\tilde{D}_{g_{*}}(w)\geq\lambda$ for some $\lambda>1$ and all $w$ in a small annulus around $S^{1}$. It follows that there is a neighborhood $V_{*}$ of $S^{1}$ which is symmetric w.r.t. $S^{1}$ such that $g_{*}^{-1}(V_{*})$ is compactly contained in $V_{*}$. If we let $V=\Psi^{-1}(V_{*}\cap\mathbb{D}^{*})\cup K$ then $V$ is a simply connected neighborhood of $K$ such that $\overline{V^{\prime}}\subset V$ where $V^{\prime}=g^{-1}(V)$.
@@
Second proof. Here we don’t use that $K$ is a continuum. For every $j\geq 0$, let $\Omega_{j}:=\cup_{i=j}^{j+n-1}g^{-i}(V_{0})$. As $V_{-n}=g^{-n}(V_{0})$ and $\overline{V_{-n}}\subset V_{0}$, then $K=K_{G}=\cap_{i>0}g^{-ni}(V_{0})$ Hence, by continuity, there exists $j_{*}>0$ such that, for every $j\geq j_{*}$, $K\subset\Omega_{j}\subset V_{0}$ and $\Omega_{j}$ shrink to $K$ as $j\to\infty$. This allows us to write, for every $j\geq j_{*}$:
$g^{-1}(\Omega_{j})=\cup_{i=j+1}^{j+n}g^{-i}(V_{0})=\cup_{i=j+1}^{j+n-1}g^{-i}(V_{0})\cup g^{-j-n}(V_{0})=\Omega_{j}\setminus g^{-j}(V_{0})\cup g^{-j-n}(V_{0}).$
Therefore, $g^{-1}(\Omega_{j})\subset\Omega_{j}$ and
$g:g^{-1}(\Omega_{j})\setminus K\to\Omega_{j}\setminus K$
@@
Let $\Omega^{*}=\Omega_{j_{*}}\setminus K$ and let $\rho_{*}$ be the hyperbolic metric of the domain $\Omega^{*}$. As $g^{-1}(\Omega^{*})\subset\Omega^{*}$ and there is no an equality, by the above and the Schwartz lemma, $g^{-1}:\Omega^{*}\to\Omega^{*}$ is a local strict contraction.
@@
Let $Y$ be the topological hull of $\overline{\Omega_{j}}$. As $g^{-1}(\overline{\Omega_{j}})\subset\overline{\Omega_{j}}$ and $g:U^{\prime}\to U$ is a proper map between simply connected domains, $g^{-1}(Y)\subset Y$. Hence, if $Y^{*}=Y\setminus K$, then $g^{-1}(Y^{*})\subset Y^{*}$ as well (remind that $K$ is completely invariant).

### v2

@@ 3. On BI maps
Let’s verify the last claim. Let $W$ be a neighborhood of $Y$. Fix $j$. By item (3), $g_{0}^{-j}(x_{0})=(g_{0}^{-1})^{j}(x_{0})$ for $x_{0}\in Y$ and all preimages are also in $Y$, hence, this holds for all $x$ in a neighborhood $W_{x_{0}}$ of $x_{0}$ (with all preimages of $x$ in neighborhoods of corresponding preimages of $x_{0}$ and in $W$). By compactness of $Y$, the union of finitely many $W_{x_{0}}$ form a neighborhood $W^{\prime}_{j}$ of $Y$. Then $W^{\prime}=\cap_{j=0}^{n-1}W^{\prime}_{j}$ is as required.
@@
Next, we are going to describe $3$ alternatives for a BI map. Let $g:U^{\prime}\to U$ be a BI map, where $U^{\prime}=\cup_{i=1}^{\infty}U_{-i}$, $U=\cup_{i=0}^{\infty}U_{-i}$. Let us fix a Jordan curve $\Gamma_{0}$ in $U_{0}$ which encloses the postcritical set $P_{g}$. Let $V_{0}$ be the Jordan domain bounded by $\Gamma_{0}$, so that $\overline{V_{0}}\subset U_{0}$.
Recall that $g_{i}^{-n}:U_{-i}\to U_{-i-n}$ is the inverse (multi-values) function to $g^{n}:U_{-i-n}\to U_{-i}$.
@@
Let, for each $n>0$, $V_{-n}=g_{0}^{-n}(V_{0})$ (which is compactly contained in $U_{-n}$) and $\Gamma_{n}=\partial V_{-n}$ (which is also a Jordan curve).
@@
Assume that (A) holds. Then, for every $\varepsilon>0$ and every $n$ big enough, the map $g^{n}:U_{-n}\to U_{0}$ has a repelling fixed point in $\varepsilon$-neighborhood of $\Gamma_{0}$.
Assume that (B) holds. Then $g^{n}:V_{-n}\to V_{n}$ is a PL map of degree $d^{n}$. Assume additionally that there is a full compact $K\subset\cap_{i\geq 0}U_{-i}$ as in Theorem 1.1, i.e., $K\supset P_{g}$ and $g^{-1}(K)=K$. Then there exist a polynomial-like restriction $g:V^{\prime}\to V$ where $V$ is a simply connected neighborhood of $K$ and $K$ is the filled Julia set of $g:V^{\prime}\to V$.
Assume (C) holds. Then there is an attracting fixed point $a\in\cap_{i\geq 0}g^{-i}(V_{0})$ such that $g^{n}(z)\to a$ for all $z\in\cup_{i\geq 0}g^{-i}(V_{0}$. Moreover, $g$ has no a completely invariant compact. Finally, if the alternative (C) holds for a sequence of $\Gamma_{0}$’s that tends to $\partial U$, then the BI map $g:U^{\prime}\to U$ is trivial: there is an attracting fixed point $a\in\cap_{i\geq 0}U_{-i}$ such that $g^{n}(z)\to a$ for all $z\in U$.
@@
(A). In this case, there exist a sequence of indices $n_{i}\to\infty$ and a sequences of points $x_{i}\in\Gamma_{n_{i}}\cap\Gamma_{0}$ as follows: $x_{i}\to b^{\prime}$, $y_{i}:=g^{n_{i}}(x_{i})\in\Gamma_{0}$, $y_{i}\to b$ and $(g^{n_{i}})^{\prime}(x_{i})\to L$ where $L\in\mathbb{C}\cup\{\infty\}$. Fix a small disk $B$ centered at $b$ which is contained in $U_{0}$ and disjoint with $P_{g}$, and, for each $i$, let $\phi_{i}$ be a branch of $g_{0}^{-n_{i}}$ which is defined in $B$ and maps $y_{i}$ to $x_{i}$. Then $(\phi_{i})^{\prime}(y_{i})\to 1/L$.
Thus $\phi_{i}:B\to\mathbb{C}$, $i=1,2,$..., is a sequence of univalent maps and is a normal family because it avoids taking values in the set $P_{g}\cup g^{-1}(P_{g})$, which must contain at least $2$ points by Definition 1.1. By this reason, the case $L=0$ is impossible.
Consider the case $L\neq\infty$, that is, $0<|L|<\infty$. Replacing $(n_{i})$ by a subsequence, in this case $(\phi_{i})$ tends uniformly on compacts in $B$ to a non-constant, hence, univalent function $\nu$ which maps $B$ to a neighborhood $B^{\prime}$ of $b^{\prime}$. As $\psi_{i}(B)\subset g_{0}^{-n_{i}}(U_{0})=U_{-n_{i}}$, then $B^{\prime}\subset U_{-n_{i}}$ for all $i$ big enough. Shrinking slightly $B^{\prime}$, we get:
@@
Let $m_{i}=n_{n+1}-n_{n_{i}}$ Passing again to a subsequence of $(n_{i})$, one can assume that $m_{i}\to\infty$. Write $g^{n_{i+1}}=g^{m_{i}}\circ g^{n_{i}}$, on $B^{\prime}$. Here, $g^{n_{i}}(B^{\prime})\subset g^{n_{i}}(U_{-n_{i+1}})\subset U_{-m_{i}}$.
Shrinking again $B^{\prime}$ if necessary, by the Claim, $g^{m_{i}}\to\operatorname{Id}$, uniformly on $\nu^{-1}(B^{\prime})$ where $\nu^{-1}(B^{\prime})\subset U_{-m_{i}}$. Concluding this argument (and shrinking $B$ to a smaller disk around $b$), we obtain that, for some branches $I_{i}:=g_{0}^{-m_{i}}$ on $B$, $I_{i}\to\operatorname{Id}$ uniformly on $B\subset U_{0}$.
Let now $D$ be any simply connected subdomain of $U_{0}\setminus P_{g}$ containing $B$. As all branches of $g_{0}^{-m_{i}}$ are well-defined in $D$ and form a normal family, passing to a subsequence, we get that $I_{i}$ (more precisely, their extensions from $B$ to $D$) converge uniformly on compacts in $D$ to $\operatorname{Id}$. As each $I_{i}$ is an inverse to $g^{m_{i}}$, it follows that, given a compactly contained in $D$ domain $F\subset D$, for all $i$ big enough, we have: $F\subset U_{-m_{i}}$ and $g^{m_{i}}\to\operatorname{Id}$ uniformly on $F$.
We apply this to two domains $D_{1},D_{2}$ which together cover the curve $\Gamma_{0}$, and to their compactly contained subdomains $F_{1}\subset D_{1}$, $F_{2}\subset D_{2}$, such that $F_{1}\cup F_{2}\supset\Gamma_{0}$. As a result, we obtain that $\Gamma_{0}\subset U_{-m_{i}}$ for all $i$ big enough and $g^{m_{i}}\to\operatorname{Id}$ uniformly on $\Gamma_{0}$. This is a contradiction by the following reason: on the one hand, as we have just shown, $\Gamma_{0}\subset U_{-m_{i}}$, also $\Gamma_{0}=\partial V_{0}$ and $g^{m_{i}}(U_{-m_{i}})=U_{0}$, hence, the sequence $g^{m_{i}}$ is well-defined on $\overline{V_{0}}$ and converges uniformly to the identity map $\operatorname{Id}$. Then $(g^{m_{i}})^{\prime}\to 1$ uniformly on compacts in $V_{0}$. On the other hand, $(g^{m_{i}})^{\prime}(c)=0$ for each $c\in C_{g}$ This contradiction excludes the case $0<|L|<\infty$.
@@
(B). $\overline{V}_{-n}\subset V_{0}$ for some fixed $n>0$. Then $G:=g^{n}:V_{-n}\to V_{0}$ is a polynomial-like map (of degree $d^{n}$). Let $K_{G}$ be the filled Julia set of $G$. Since $g^{-1}(K)=K$ and using Lemma 3.1, $G^{-1}(K)=g_{0}^{-n}(K)=K$. As $G:V_{-n}\to V_{0}$ is conjugate to a polynomial and $K\geq 2$, it follows that $K=K_{G}$. Moreover, as $P_{G}=P_{g}\subset V_{0}$, hence, $K=K_{G}$ is connected.
@@
It remains to show $g$ itself has a PL restriction, i.e., there exist a simply connected neighborhood $V$ of $K$, such that $V^{\prime}:=g^{-1}(V)$ is compactly contained in $V$. If $n=1$, there is nothing to prove (set $V=V_{0}$).
@@
So let $n\geq 2$. We present two proofs: the first one uses heavily that $K$ is a continuum while the second one is "continuum free" and can be easily adapted to the set up of Theorem 1.3 where $K$ is disconnected.
@@
First proof. The following argument is well-known. Let $\Psi:\mathbb{C}\setminus K\to\mathbb{D}^{*}:=\{|w|>1\}$ be the conformal isomorphism such that $\Psi(\infty)=\infty$. Below, $\hat{A}=\Psi(A\setminus K)$ for a neighborhood $A$ of $K$, and $A_{*}$ is the union of $\hat{A}$ with its symmetric w.r.t. $S^{1}$ completed by $S^{1}$. Similar notations are used for maps. Replace $g:U_{-1}\setminus K\to U_{0}\setminus K$ by $\hat{g}:=\Psi\circ g\circ\Psi^{-1}:\hat{U}_{-1}\to\hat{U}_{0}$. By Schwartz’s reflexion principle, $\hat{g}$ extends to a holomorphic map $g_{*}:(U_{-1})_{*}\to(U_{0})_{*}$. The latter map has no critical points as $g_{*}^{-1}(S^{1})=S^{1}$. Let $\hat{G}=\Psi\circ g\circ\Psi^{-1}:\hat{U}_{-m}\to\hat{U}_{0}$ and $G_{*}:(U_{-m})_{*}\to(U_{0})_{*}$. As this map is a covering and $\overline{(U_{-m})_{*}}\subset(U_{0})_{*}$, then $G_{*}$ expands the hyperbolic metric $\rho$ of the domain $(U_{0})_{*}$, i.e, $D_{G}(w):=\rho\circ G(w)|G^{\prime}(w)|/\rho(w)>1$ for $w\in(U_{-m})_{*}$. Now note that $G_{*}=g_{*}^{m}$ in some neighborhood of $S^{1}$. Define a new metric $\tilde{\rho}=\sum_{i=0}^{m-1}\rho\circ g_{*}^{i}|(g_{*}^{i})^{\prime}|$. Then $\tilde{D}_{g_{*}}(w):=\frac{\tilde{\rho}\circ g_{*}(w)|g_{*}^{\prime}(w)|}{\tilde{\rho}(w)}=1+\frac{\rho(w)}{\tilde{\rho}(w)}\{D_{G}(w)-1\}>1$ for all $w$ near $S^{1}$. By the compactness, $\tilde{D}_{g_{*}}(w)\geq\lambda$ for some $\lambda>1$ and all $w$ in a small annulus around $S^{1}$. It follows that there is a neighborhood $V_{*}$ of $S^{1}$ which is symmetric w.r.t. $S^{1}$ such that $g_{*}^{-1}(V_{*})$ is compactly contained in $V_{*}$. Then we let $V=\Psi^{-1}(V_{*}\cap\mathbb{D}^{*})\cup K$.
Second proof. Here we don’t use that $K$ is a continuum.
As $V_{-n}=g_{0}^{-n}(V_{0})$ and $\overline{V_{-n}}\subset V_{0}$, then the iterates $(g_{0}^{-n})^{j}$ ($j>0$) of $g_{0}^{-n}$ are all well-defined on $V_{0}$ and $\tilde{V}_{j}:=(g_{0}^{-n})^{j}(V_{0})$, $j=0,1,$..., form a decreasing sequence of Jordan domains (with $\tilde{V}_{0}=V_{0}$)such that
(3.1) $K=K_{G}=\cap_{j>0}(\tilde{V}_{j}).$
Note that $\tilde{V}_{j+1}=g_{0}^{-n}(\tilde{V}_{j})$ and $\overline{\tilde{V}_{j+1}}\subset\tilde{V}_{j}$.
For every $j\geq 0$, let $\Omega_{j}:=\cup_{i=0}^{n-1}g_{0}^{-i}(\tilde{V}_{j})$.
Apply the last claim of Lemma 3.1 to a neighborhood $W=V_{0}$ of $K$ and find a neighborhood $W^{\prime}$ of $K$ such that
(3.2) $g_{0}^{-i}=(g_{0}^{-1})^{i}\mbox{ for }i=0,1,...,n.$
@@
By (3.1), there exists $j_{*}>0$ such that, for every $j\geq j_{*}$, $K\subset\Omega_{j}\subset W^{\prime}$ and $\Omega_{j}$ shrink to $K$ as $j\to\infty$.
This allows us to write, using (3.2), for every $j\geq j_{*}$:
$g_{0}^{-1}(\Omega_{j})=g_{0}^{-1}(\cup_{i=0}^{n-1}(g_{0}^{-1})^{i}(\tilde{V}_{j}))=\cup_{i=1}^{n}(g_{0}^{-1})^{i}(\tilde{V}_{j})=$
$\Omega_{j}\cup[(g_{0}^{-1})^{n}(\tilde{V}_{j})\setminus\tilde{V}_{j})]=\Omega_{j}\cup[(g_{0})^{-n}(\tilde{V}_{j})\setminus\tilde{V}_{j})]=$
$\Omega_{j}\cup\tilde{V}_{j+1}\setminus\tilde{V}_{j}\subset\Omega_{j}.$
Thus $g_{0}^{-1}(\Omega_{j})\subset\Omega_{j}$ and
$g:g_{0}^{-1}(\Omega_{j})\setminus K\to\Omega_{j}\setminus K$
@@
Let $\Omega^{*}=\Omega_{j_{*}}\setminus K$ and let $\rho_{*}$ be the hyperbolic metric of the domain $\Omega^{*}$. As $g_{0}^{-1}(\Omega^{*})\subset\Omega^{*}$ and there is no an equality, by the above and the Schwartz lemma, every local branch of $g_{0}^{-1}:\Omega^{*}\to\Omega^{*}$ is a strict contraction.
@@
Let $Y$ be the topological hull of $\overline{\Omega_{j}}$. As $g_{0}^{-1}(\overline{\Omega_{j}})\subset\overline{\Omega_{j}}$ and $g:U_{-1}\to U_{0}$ is a proper map between simply connected domains, $g_{0}^{-1}(Y)\subset Y$. Hence, if $Y^{*}=Y\setminus K$, then $g_{0}^{-1}(Y^{*})\subset Y^{*}$ as well (remind that $K$ is completely invariant).